# Azumaya algebras and Barr’s Theorem

### v3

@@ 2 Some lemmas about Henselian rings
Let $u$ be a vector column of generators of $M$: then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have (IP)u=v(I-P)u=v and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
Corollary 2.1.
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$. Then $u$ is invertible in $A$.
Proof.
We have $A=uA+\mathsf{m}A$ and we can apply the previous Lemma. ∎
Lemma 2.3 ([ALP, Prop. 5.8]).
Let $P$ in $R[X]$ be a monic polynomial in $R[X]$ and assume that we have a decomposition $P=fg$ mod. $\mathsf{m}R[X]$ with $f$ and $g$ monic polynomials coprime mod. $\mathsf{m}$. Then we can find a decomposition $P=FG$ in $R[X]$ with $F=f$ and $G=g$ mod. $\mathsf{m}$.
Proof.
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that (f)=(e),(g)=(1e)(f)=(e),~(g)=(1-e) mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that (1e)F(x)=0(1-e)F(x)=0 in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
Theorem 2.1 ([ALP, Thm. 5.9]).
If $A$ is an $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathsf{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathsf{m}A$.
Proof.
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0)$, $h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that 1a1-a is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~Q=(1-X)^{m},~H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that X(1X)U+HV=Res(X(1X),H)=1X(1-X)U+HV=Res(X(1-X),H)=1 [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and Q(a)=1aQ(a)=1-a mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
@@ 3 Azumaya algebras
    AA is projective of finite type;
@@
    the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
@@
One can equivalently describe the algebra structure by a bilinear map $\mu:A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors. Thus all subsequent arguments can be phrased without explicit tensor notation.
Over a finite free $k$-module, to give an algebra structure is represented concretely by a multiplication table. A module $A$, projective of finite type over $k$, is given concretely by a projection matrix over $k$ [LQ], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$). We can then compute [LQ, X, Theorem 1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A_{s_{i}}$ is free over $k_{s_{i}}$, and the second condition for being an Azumaya algebra becomes then that some determinants $\delta_{1},\dots,\delta_{l}$ are invertible [^5] After localisation to $s_{j}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$.. From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a $k$-algebra $k\rightarrow S$ then the algebra $A_{S}$ over $S$, with the same multiplication in $S$, is also an Azumaya algebra.
We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
@@
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree with root $k$ using the following, where at all leaves all scalar extensions $A_{S}$ are matrix algebras:
@@
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
We prove first that if we have such a tree then the algebra $A$ is Azumaya. If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension. (As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section. (This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented $k$-algebra, is, in the internal logic of this topos, a local strictly Henselian ring. The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is an Azumaya algebra, given by the same multiplication table as the one of $A$. If we can prove in intuitionistic logic that an Azumaya algebra over a strictly Henselian ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree of $k$ which gives Theorem 3.1.
Internally in the classifying topos of stricly Henselian $k$-algebras, the object $R$ is local strictly Henselian and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable intuitionistically. Such a proof yields externally a finite covering tree of $k$ satisfying the conditions above.
@@
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and separably closed $k$-algebra, then $A$ is a matrix algebra.
@@
We know that $R$ is Henselian and that $R/\mathsf{m}$ is separably closed by Theorem 1.1.
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism. ∎
The next proof uses Barr’s Theorem, similar to the way it is used in [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
@@
If $A$ is an Azumaya algebra over $R$, which is a local and separably closed $k$-algebra, then $A$ is a matrix algebra.
@@
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebras.
@@
We can use such a tree to give arguments by “tree induction”. We give some simple examples.
@@
This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $R$. We prove this by “tree induction” using Theorem 3.1. The statement holds on the leaves (it holds for a matrix algebra) and descends to the root. ∎
@@
If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
@@
Let $A=End_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$. Let $\psi$ be an automorphism of $A$. If $p_{ij}=\psi(e_{ij})$, we have that the $p_{i}=p_{ii}$ form an FSOI and $p_{ij}p_{kl}=\delta_{jk}p_{il}$. Since $R$ is local, each $V_{i}=p_{i}(R^{n})$ is free. Furthermore $R^{n}$ is a direct sum of the $V_{i}$’s; the $V_{i}$’s are pairwise isomorphic since $p_{ji}$ defines an isomorphism between $V_{i}$ and $V_{j}$. So they are all free of rank $1$. Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$. If $a(v_{i})=w_{i}$ we have $\psi(x)=axa^{-1}$. ∎
For the next result, we use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]). For completeness, we recall the statement. Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finite projective if, and only if, it is finitely presented and flat.
Theorem 3.2 (Descent).
Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module. If $M\otimes_{A}B$ is finite projective, then so is $M$. If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M_{s_{i}}$ is finite projective over $A_{s_{i}}$ then so is $M$.
@@
$M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ and $N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
@@
We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎
@@ 4 Theory of algebraically closed local rings
$\exists_{x}~P(x)=0$
for any monic nonconstant polynomial $P$.
Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s PhD thesis [BT].
@@
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and algebraically closed $k$-algebra, then $A$ is a matrix algebra.
@@
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
@@
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic nonconstant.
@@ 5 Conclusion
This paper is a beginning of the theory of Azumaya algebras in a constructive setting. The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $GL_{n}(R)$ for the étale topology.

### v4

@@ 2 Some lemmas about Henselian rings
Write $\varphi(x)=x^{2}-x$ and consider the finitely generated algebra $B=R[a]$. Since $\varphi(a)=a^{2}-a$ is in $\mathfrak{m}A$, the constant term of its characteristic polynomial is in $\mathfrak{m}A\cap R=\mathfrak{m}$, so that we have $N$ such that $\varphi(a)^{N}$ is in $\mathfrak{m}B$. We can then consider the following Newton quadratic iteration of $a$ inside $B$ by defining inductively $u_{0}=a$ and $u_{n+1}=u_{n}-(2u_{n}-1)\varphi(u_{n})$ (compare [LQ, Corollary III.10.4]). One checks that $u_{n}=a$ mod. $\mathfrak{m}A$ and that $\varphi(u_{n})$ is in $\varphi(u_{n-1})^{2}B$, so that it is in $\mathfrak{m}B$ for $n$ big enough, i.e. $u_{n}$ is idempotent mod. $\mathfrak{m}B$. We can now use Theorem 2.1 for the commutative $R$-algebra $B$ to lift $u_{n}$ to an idempotent $e$ in $B$ such that $e=u_{n}$ mod. $\mathfrak{m}B$. We also have $e=a$ mod. $\mathfrak{m}A$. ∎
@@ 3 Azumaya algebras
    AA is finitely generated projective;
@@
    the canonical map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is an isomorphism.
@@
In this section $k$ will denote a commutative base ring, and $R$ will denote the “generic” separably closed commutative $k$-algebra, model of the coherent theory $T_{1}$ described in the first section.
One can equivalently describe the algebra structure by a bilinear map $\mu\colon A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors. Thus all subsequent arguments can be phrased without explicit tensor notation.
Over a finitely generated free $k$-module, an algebra structure is given concretely by a multiplication table. A module $A$, finitely generated projective over $k$, is given concretely by a projection matrix over $k$ [LQ, Chap. I], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$). We can then compute [LQ, Theorem X-1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A[1/s_{i}]$ is free over $k[1/s_{i}]$, and the second condition for being an Azumaya algebra becomes then that $l$ determinants are invertible. [^5] After localisation to $s_{i}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$. From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a commutative $k$-algebra $k\rightarrow S$, then the algebra $A_{S}$ over $S$, with the same multiplication table, is also an Azumaya algebra.
We have the following characterisation, which generalises the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
@@
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite covering tree rooted at $k$ using the following, where at every leave the scalar extension $A_{S}$ is a matrix algebra:
@@
    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic and unramifiable.
We shall use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]). For completeness, we recall the statement. Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finitely generated projective if, and only if, it is finitely presented and flat.
Theorem 3.2 (Descent).
Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module. If $M\otimes_{A}B$ is finitely generated projective, then so is $M$. If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M[1/s_{i}]$ is finitely presented and projective over $A[1/s_{i}]$ then so is $M$.
We prove first that if we have such a covering tree then the algebra $A$ is Azumaya.
If we have such a covering tree rooted at $k$, notice that each extension $S\rightarrow S[X]/(f)$ for $f$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/s_{i}]$ is also faithfully flat if $1=(s_{1},\dots,s_{n})$. It follows then that $A$ is finitely generated projective by Theorem 3.2. The map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension. (As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section. (This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented commutative $k$-algebra, is, in the internal logic of this topos, a separably closed local ring. The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is the Azumaya algebra given by the same multiplication table as the one for $A$. If we can prove in intuitionistic logic that an Azumaya algebra over a separably closed ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree rooted at $k$ which yields Theorem 3.1.
Internally in the classifying topos of separably closed local commutative $k$-algebras, the object $R$ is separably closed local and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable in intuitionistic logic. Such a proof yields externally a finite covering tree rooted at $k$ satisfying the conditions above.
@@
If $A$ is an Azumaya algebra over $R$, which is a residually discrete separably closed local commutative $k$-algebra, then $A$ is a matrix algebra over $R$.
@@
We know that $R$ is Henselian and that $R/\mathfrak{m}$ is separably closed by Theorem 1.1.
We simplify (and correct slightly) the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathfrak{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathfrak{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathfrak{m}A$. We lift the idempotent $e_{11}$ mod. $\mathfrak{m}A$ to some idempotent $e$ in $A$ using Corollary 2.1.
The $R$-module AeAe is finitely generated projective and hence free since $R$ is a local ring. The map $A\rightarrow\operatorname{End}_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism. ∎
The next proof uses Barr’s Theorem in a way similar to [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
@@
If $A$ is an Azumaya algebra over $R$, which is a separably closed local commutative $k$-algebra, then $A$ is a matrix algebra over $R$.
@@
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of separably closed local commutative $k$-algebras.
@@
We can use such a covering tree to give arguments by “tree induction”. We give some simple examples.
@@
This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $k$. We prove this by tree induction using Theorem 3.1. The statement holds on the leaves (it holds for a matrix algebra) and descends to the root. ∎
@@
If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Definition X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $\operatorname{GL}_{n}(A)$.
@@
Let $A=\operatorname{End}_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$. Let $\psi$ be an automorphism of $A$. If $p_{ij}=\psi(e_{ij})$, we have that the $p_{i}=p_{ii}$ form an FSOI and $p_{ij}p_{kl}=\delta_{jk}p_{il}$. Since $R$ is local, each $V_{i}=p_{i}(R^{n})$ is free. Furthermore $R^{n}$ is a direct sum of the $V_{i}$’s; the $V_{i}$’s are pairwise isomorphic since $p_{ji}$ defines an isomorphism between $V_{i}$ and $V_{j}$. So they are all free of rank $1$. Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$. If $a(v_{i})=w_{i}$ then we have $\psi(x)=axa^{-1}$. ∎
@@
$M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}\quad\text{and}\quad N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
@@
We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow\operatorname{Hom}(N,k),~x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this statement also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎
@@ 4 Theory of algebraically closed local rings
$\exists_{x}~f(x)=0$
for any monic nonconstant polynomial $f$.
Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s Ph.D. thesis [BT].
@@
If $A$ is an Azumaya algebra over $R$, which is a residually discrete algebraically closed local commutative $k$-algebra, then $A$ is a matrix algebra.
@@
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite covering tree rooted at $k$ using the following, where all leaves are matrix algebras:
@@
    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic nonconstant.
@@ 5 Conclusion
This paper is a beginning of the theory of Azumaya algebras in a constructive setting. The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $\operatorname{GL}_{n}(k)$ for the étale topology.